# What Is the Geometry of the Alignment Tax?

### v1

@@ 10 Limitations
The scaling law (Theorem 5.4) assumes that incidental overlaps arise from feature packing with bounded coherence. Real models learn structured representations, not random packings. The qualitative prediction (some tasks have reducible tax, others do not) is robust to model details, but the specific rate $O(m^{\prime}/d)$ depends on the packing assumption. We frame the scaling law as a prediction conditioned on the packing hypothesis, with an experimental protocol for testing it as future work.
We model safety as a single direction or low-dimensional subspace. Fine-grained safety distinctions (is this chemistry question about homework or synthesis?) may require higher-dimensional representations. The formalization accommodates arbitrary $k=\dim(S)$ but the scaling law’s tractability benefits from $k$ being small.
Finally, the analysis is about average-case alignment (perturbations $\delta$ that work across the data distribution). Adversarial robustness (ensuring safety holds for worst-case inputs) involves the input-dependence of the Jacobian $J(x)$ and is not captured by the population-average analysis. The alignment tax for adversarial safety is likely higher, and its geometry is likely more complex, than what the current work describes.
11 Conclusion

### v2

@@ 8 Conclusion
The scaling law (Theorem 13) assumes that incidental overlaps arise from feature packing with bounded coherence. Real models learn structured representations, not random packings. The qualitative prediction that some tasks have reducible tax and others do not is robust to model details, but the specific rate $O(m^{\prime}/d)$ depends on the packing assumption. We frame the scaling law as a prediction conditioned on the packing hypothesis, with an experimental protocol for testing it as future work.
We model safety as a single direction or low-dimensional subspace. Fine-grained safety distinctions like whether this chemistry question about homework or synthesis may require higher-dimensional representations. The formalization accommodates arbitrary $k=\dim(S)$ but the scaling law’s tractability benefits from $k$ being small.
Finally, the analysis is about average-case alignment, namely perturbations $\delta$ that work across the data distribution. Adversarial robustness (ensuring safety holds for worst-case inputs) involves the input-dependence of the Jacobian $J(x)$ and is not captured by the population-average analysis. The alignment tax for adversarial safety is likely higher, and its geometry is likely more complex, than what the current work describes.